A student takes an exam containing 18 true or false questions. If the student guesses, what is the probability that he will get exactly 15 questions right? Round your answer to four decimal places.
Answer #1

Here, n = 18, p = 0.5, (1 - p) = 0.5 and x = 15
As per binomial distribution formula P(X = x) = nCx * p^x * (1 - p)^(n - x)

We need to calculate P(X = 15)


P(x =15) = 18C15 * 0.5^15 *(1-0.5)^3
= 0.0031